OpenStudy (anonymous):

match n = -32 n = -0.2 n = -5 n = -3.2 with n/8 + 1/2 = -3 1/2 -1.4n + 2.1 = 6.58 5.7n + 4.1 = 2.96 4n - 2 3/4 = -22 3/4

OpenStudy (gorv):

n/8=-31/2-1/2=-16 n=-16*8=-128